OpenStudy (anonymous):

There were 150 people at the Junior-Senior dance Sasha and her friends attended. Junior tickets were $2.00 each and Senior tickets were $3.50 each. The total receipts for the dance were $405. How many Juniors bought tickets?

OpenStudy (anonymous):

o easy u set up two equations for this one.

OpenStudy (anonymous):

u wanna know how?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

ok so for the first equation ur going to set it up as j+s=150 which is j=juniors, s=seniors and 150 is equal to the total # of ppl that attended. ok, so the next equation would be 2j+3.50s=450 which again juniors , seniors and 450 is the total receipt for the dance..... get it so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok good! do u know how to cancel out variables?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okay so then u would simply cancel out juniors and solve for s/seniors! :)

OpenStudy (anonymous):

2(150-s)+3.50s=450 so that will be the problem?

OpenStudy (anonymous):

let me see........

OpenStudy (anonymous):

it this what u got after getting rid of j?

OpenStudy (anonymous):

well yea.. didnt you get that?

OpenStudy (anonymous):

ahhhhhhh no welll..... when i canceled out the j u got.... well ok this is before so, i had -2j-2s=-300 and 2j+3.50s=450 then i canceled the j and i got 1.5s=150

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

100?

OpenStudy (anonymous):

yeah u got it! :) good job!!!!!!!!!

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

no problem. :)